OpenStudy (anonymous):

A culture of bacteria contains 150 bacteria at time zero. Find the number of bacteria after 5 hours if their number doubles every hour. I know this problem is simple but I'm wondering if someone could solve this with a formula that i can use on more difficult problems? thanks!

OpenStudy (anonymous):

So here you can use an exponential equation of \[y=P(r)^x\] Where P is the initial number, r is the rate of change, and x is the time So P is 150 because that's what you start with, and r is 2 because you double it You would substitute 5 for x and solve

OpenStudy (anonymous):

ah ok thank you so much!!

OpenStudy (jdoe0001):

if we start off, with, say... some value and then we double it on the next so, say we start with 7 so \(\begin{array}{llll} value&doubled \\\hline\\ 7&7\cdot 2\implies 7(2^1)\\ (7\cdot 2)&(7\cdot 2)\cdot 2\implies 7(2^2)\\ (7\cdot 2\cdot 2)&(7\cdot 2\cdot 2)\cdot 2\implies 7(2^3)\\ ...&.... \end{array}\) notice, what the equation would be?

OpenStudy (jdoe0001):

it just so happens that in your case, is 150, not 7 but if you set 2 with an exponential of 0, you'll see what that'd give you

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

thanks i really appreciate that